LootAtTheTime:

Evaluate the imaginary number i^3

dude:

-i

LootAtTheTime:

Ok thats what I got dude but can you explain it

LootAtTheTime:

Please?

LootAtTheTime:

I’m trying to solve another and an explanation o this one would help a lot

Marziman:

Essentially, i^3 would be equivelant to i^2 * i and so on and so forth.

dude:

So every power that `i`is raised to follows the set of numbers \(i, -1,-i, 1\) \(i^1=i\\ i^2=-1\\ i^3=-i\\ i^4=1\)

Marziman:

thereby making i^2 * i = -1 * i leaving you with -i

dude:

As you continue on with the exponents the cycle continues \(i^5=i\\ i^6=-1 \\ i^7=-i\\ i^8=1\)

LootAtTheTime:

That’s confusing but I think I get it thanks guys!

TheSmartOne:

If we start from the basics, do you understand that \(\Large i = \sqrt{-1}\)

Marziman:

average math noobie: "NO TSO THAT'S WRoOnGG IT CAN'T = -1!!!!111"

Marziman:

@TheSmartOne